OpenStudy (anonymous):

Tori's 8-ounce glass of soymilk has 6 grams of protein. Brady's 16-ounce glass has 12 grams of protein. What is ratio of Tori's drink size to Brady's drink size? Multiple choice: 1:1, 4:3, 16:4, 3:4.

OpenStudy (anonymous):

@king can u help me plzz

OpenStudy (king):

thinkin!

OpenStudy (king):

i think its 1:1...do u hav the answer....is my answer rite?

OpenStudy (anonymous):

yes tthxx

OpenStudy (king):

see in Tori's glass protein is in the ratio 6/8=3/4 and in Brady's glass protein is in ratio 12/16=3/4 so the ratio of Tori and Brady will be 3/4 by 3/4=1/1=1:1